OpenStudy (anonymous):

If f(4) = 8, g(4) = 2, f(-4) = 1, g(-4) = 3, f'(4) = 10, g'(4) = 12, f'(-4) = 6, and g'( -4) = -2, find H'(4) where H(x) = g( f(x)- 3x).

OpenStudy (anonymous):

how do i solve this?

OpenStudy (anonymous):

h= g (f(x)-3x) h'(x) = ( f '(x) -3 ) g ' ( f(x) -3x) , by chain rule

OpenStudy (anonymous):

now sub the value h'(4) = ( f'(4) -3) g ' ( f(4) -12)

OpenStudy (anonymous):

H'(x) = g'(f(x) - 3x) * (f'(x) - 3) H'(4) = g'(f(4) - 3*4) * (f'(4) - 3) = g'(8 - 12) * (10 - 3) = g'(-4) * 7 = -2*7 = -14

OpenStudy (anonymous):

but f(4) =8 , and f'(4) = 10

OpenStudy (anonymous):

etc ^

OpenStudy (anonymous):

thank you!!!!!:))

OpenStudy (anonymous):

chain rule confuses me so much